### v1

@@ 5 Conclusion
In this paper, we propose VLM guidance-based semi-supervised CD method DiffMatch, aiming to explore the effectiveness of VLM in CD tasks. To sufficiently and rationally utilize the pseudo labels generated by VLM, we introduce five components in DiffMatch, which not only provide abundant and reliable pseudo labels for unlabeled data, but also explicitly decouple the CD process. We conduct extensive experiments on two RSCD datasets and the results demonstrate the superiority of DiffMatch. In addition, our proposed CEG strategy makes a performance leap in un-supervised CD. Further, DiffMatch is also applicable to a wide range of CD tasks in other scenarios.
@@
The limitation of DiffMatch is that generating pseudo labels is time-consuming, which requires extra training cost, or pre-generating the required VLM pseudo labels. The merit of this research is that it demonstrates the possibilities and potential of VLMs in semi-supervised and un-supervised CD tasks, which is a direct or indirect path to realize a universal CD model in the future.
@@ Acknowledgements
The authors would like to thank Xiaoliang Tan and Guanzhou Chen for some experimental data on un-supervised CD methods (Table 5).
[…]

### v2

@@ V Conclusion
In this paper, we propose a VLM guidance-based semi-supervised CD method DiffMatch, aiming to explore the effectiveness of VLM in CD tasks. To sufficiently and rationally utilize the pseudo labels generated by VLM, we introduce five components in DiffMatch, which not only provide abundant and reliable pseudo labels for unlabeled data, but also explicitly decouple the CD process. We conduct extensive experiments on two RSCD datasets and the results demonstrate the superiority of DiffMatch. In addition, our proposed CEG strategy makes a performance leap in un-supervised CD. Further, DiffMatch is also applicable to a wide range of CD tasks in other scenarios.
@@
There are two main limitations of DiffMatch: 1) Although we propose the mixed CEG strategy to obtain more reliable pseudo labels, it is inevitable that the incorrect pseudo labels still exist due to the imperfect VLMs. 2) The generation of pseudo labels is time-consuming, which requires extra training cost, or pre-generating the required VLM pseudo labels. The merit of this research is that it demonstrates the possibilities and potential of VLMs in semi-supervised and un-supervised CD tasks, which is a direct or indirect path to realize a universal CD model in the future.
@@ Acknowledgements
The authors would like to thank Xiaoliang Tan and Guanzhou Chen for some experimental data on un-supervised CD methods (Table V).